# Unveiling Group-Specific Distributed Concept Drift: A Fairness Imperative in Federated Learning

### v3

@@ Abstract
In the evolving field of machine learning, ensuring fairness has become a critical concern, prompting the development of algorithms designed to mitigate bias in decision-making processes. However, achieving fairness in the presence of group-specific concept drift remains an unexplored frontier, and our research represents pioneering efforts in this regard. Group-specific concept drift refers to situations where one group experiences concept drift over time while another does not, leading to a decrease in fairness even if accuracy remains fairly stable. Within the framework of Federated Learning, where clients collaboratively train models, its distributed nature further amplifies these challenges since each client can experience group-specific concept drift independently while still sharing the same underlying concept, creating a complex and dynamic environment for maintaining fairness. The most significant contribution of our research is the formalization and introduction of the problem of group-specific concept drift and its distributed counterpart, shedding light on its critical importance in the realm of fairness. Additionally, leveraging insights from prior research, we adapt an existing distributed concept drift adaptation algorithm to tackle group-specific distributed concept drift which uses a multi-model approach, a local group-specific drift detection mechanism, and continuous clustering of models over time. The findings from our experiments highlight the importance of addressing group-specific concept drift and its distributed counterpart to advance fairness in machine learning.
@@ I Introduction
Fairness is a fundamental concern in machine learning, aiming to prevent discrimination in algorithmic decision-making processes with respect to sensitive attributes, such as race or gender [1, 2]. However, achieving fairness becomes even more challenging in the presence of concept drift, which refers to the phenomenon where the underlying data distribution changes over time. In this paper, we first introduce and formalize the concept of group-specific concept drift, which focuses on scenarios where one group, characterized by a sensitive attribute, experiences changes in the data distribution while another group remains unaffected.
@@
To illustrate the importance of the problem, consider an example of group-specific concept drift in the context of loan lending. Suppose a financial institution operates in a region where historically marginalized communities have limited access to financial services. Over time, societal changes, government initiatives, and increased awareness promote financial inclusion, leading to an increase in loan applications from these previously underserved communities. In this evolving environment, group-specific concept drift occurs as the data distribution for loan applicants from marginalized communities changes. Traditional lending models may not adequately capture the creditworthiness and financial profiles of these communities due to the historical underrepresentation and lack of relevant data. Consequently, there is a risk of perpetuating bias in loan approval decisions. Addressing group-specific concept drift is crucial to ensure fair access to loans for individuals from historically marginalized communities.
Federated Learning (FL), a decentralized approach to machine learning, provides a unique setting for studying fairness under group-specific distributed concept drift. In this framework, multiple clients collaboratively train models using their local data, with a central server aggregating their updates to create a global model, preserving client data privacy and sharing only model updates. [3]. However, the distributed nature of FL, combined with the presence of group-specific concept drift, poses new challenges that have not been addressed in the context of fairness. In this context, group-specific concept drift could manifest in specific clients at different times, creating challenges in maintaining fairness and accuracy. As such, in this work, we also introduce and formalize the concept of group-specific distributed concept drift in distributed environments.
To illustrate the importance of the problem of group-specific distributed concept drift, consider several examples across different domains:
@@
Example 1: Going back to the previous example of loan lending, now in a FL setting, consider a scenario where clients representing different regions or demographic groups experience varying rates of change in the data distribution of loan applicants. For instance, a financial institution operating in a diverse geographic landscape may have multiple branches, each serving a unique community with its distinct characteristics. In this FL framework, if one branch undergoes changes in the data distribution for loan applicants from marginalized communities, while another branch remains unaffected, the result is a manifestation of group-specific distributed concept drift. The decentralized nature of FL implies that these shifts in data distribution occur independently across different clients at various times, introducing a layer of complexity in maintaining fairness and accuracy. As the financial landscape evolves and societal changes unfold, addressing group-specific distributed concept drift becomes imperative to uphold fairness and prevent perpetuation of biases in loan approval decisions for diverse and historically underserved communities.
Example 2: Consider an international online fashion retailer using FL to tailor product recommendations for users across different countries. The retailer’s user base encompasses a diverse array of demographic groups, each with distinct fashion inclinations shaped by cultural norms, climate, and socioeconomic factors. Envision a scenario where a particular fashion trend emerges in Country A (i.e. client A) amongst women (i.e. sensitive group - gender) due to cultural phenomena or seasonal shifts. However, this same trend might not resonate or even be visible in Country B (i.e. client B) due to differing cultural norms or consumer preferences. In some instances, what is considered fashionable in Country A might only gain traction among women in Country B after a significant delay, or it might not find relevance at all due to cultural disparities. Consequently, the online fashion retailer encounters group-specific distributed concept drift, wherein the data distribution of user preferences and purchasing behaviors diverges across different countries and within specific demographic segments. While women in Country A may eagerly adopt the latest trend, other regions such as Country B might exhibit different fashion preferences or remain indifferent to the trend altogether. This concept drift presents notable challenges for the retailer’s FL system. Failing to address this group-specific distributed concept drift could result in suboptimal recommendations, diminished user engagement, and the risk of biased or discriminatory outcomes, particularly if certain demographic segments are underrepresented in the recommendation process.
Example 3: Consider a healthcare system serving patients from diverse ages, using FL to develop predictive models for disease diagnosis and treatment recommendations. However, the healthcare landscape is dynamic, with demographic groups experiencing fluctuations in health behaviors, healthcare access, and disease prevalence. Consider a scenario where certain hospitals, operating as individual clients within the FL framework, introduce new drugs or treatment protocols. These interventions aim to address specific health conditions prevalent within their patient populations. However, the introduction of these drugs may inadvertently lead side effects within these hospitals’ patient populations. For instance, the administration of certain drugs for managing a disease may lead to side effects among older patients (i.e. sensitive group - age). Consequently, the data distribution of disease patterns within these hospitals undergoes a noticeable shift, primarily affecting the older demographic. In contrast, younger patients within these hospitals who are less susceptible to developing these side effects as a result of the drug, may not experience the same level of change in disease manifestations. This distributed concept drift primarily affects the older demographic, underscoring the importance of adapting predictive models to evolving data distributions. Addressing this group-specific distributed concept drift is essential to ensure accurate and equitable healthcare outcomes across diverse patient populations, thereby minimizing disparities in healthcare delivery.
@@
We pioneer the introduction and formalization of the concept of group-specific concept drift and group-specific distributed concept drift, shedding light on previously unexplored aspects within the fairness landscape
@@
The remainder of this work is structured as follows: Section 2 provides some background and related work on the subject, Section 3 describes the problem statement, Section 4 presents the proposed approach, Section 5 describes the methodology and experimental design, Section 6 presents the results and discussion, Section 7 presents the conclusions and possible future research.
@@ II Related Work
II-A Fairness
Fairness-aware machine learning has gained significant attention, and algorithms designed to promote fairness can be categorized into three groups based on the stage at which they are implemented: pre-processing, in-processing, and post-processing [1, 2, 4, 5]. The algorithm proposed in this study belongs to the in-processing category. Fairness metrics can be broadly classified into two main groups: group fairness and individual fairness. Group fairness emphasizes that the outcomes of an algorithm do not disproportionately favor or harm any particular group based on characteristics such as race, gender, ethnicity, religion, or socioeconomic status. On the other hand, individual fairness aims to ensure similar predictions for individuals who share similar characteristics [1, 4]. This research specifically focuses on optimizing group fairness.
@@ II-B Fair Federated Learning
In the realm of FL, there has been a growing focus on achieving fairness [6, 7, 8, 9, 10, 11]. These methods aim to mitigate bias while preserving privacy in a decentralized environment using techniques such as fair global model aggregation techniques [6, 7, 8, 9] and local debiasing methodologies at each client [10, 7].
These approaches, designed for static FL, lack mechanisms to detect and adapt to changing data distributions, rendering them unfit for addressing fairness under group-specific distributed concept drift. In such scenarios, characteristics or patterns observed in one concept may not be applicable to others, intensifying fairness challenges in FL environments. Hence, using a multi-model approach offers a promising solution to address the complexities of concept drift while preserving fairness across diverse data distributions.
@@ II-C Federated Learning under Concept Drift
Concept drift has been extensively studied in the centralized setting, and readers can refer to the surveys [12, 13] for further details. However, applying centralized algorithms to FL may not be well-suited for handling distributed concept drifts with heterogeneous data distributions across time and clients [14, 15, 16, 17]. In the context of FL, research on concept drift is still in its early stages and most works do not account for the distributed nature of concept drift in their algorithms [14].
In their recent work, Jothimurugesan et al. [14] introduce a novel approach to tackle the challenge of distributed concept drift in FL. They treat drift adaptation as a dynamic clustering problem, emphasizing the limitations of single global models when dealing with distributed drifts. To address this issue, the authors propose a multi-model approach that leverages local drift detection based on the overall loss, offering a more robust solution to adapt to evolving data distributions over time. However, it is important to note that this algorithm, referred to as FedDrift, focuses on the global loss and does not consider group-specific losses. In the context of our work, where we specifically address group-specific distributed concept drift, FedDrift’s reliance on global loss renders it ill-suited for our scenario. Group-specific concept drift may not be effectively detected by FedDrift, as it does not account for group-specific loss variations in local models that are crucial for capturing changes specific to certain groups. This limitation underscores the need for a more nuanced approach to address fairness concerns under group-specific distributed concept drift, which we elaborate on in the subsequent sections.
@@ II-D Limitations and Research Context
While certain studies have delved into fairness under concept drift in centralized streaming environments [18, 19, 20, 21], it is important to recognize their limitations when applied to fairness under concept drift in FL. These fairness-aware algorithms may not be well-suited for FL due to several reasons. Firstly, they typically require direct access to sensitive attributes, which is restricted in FL where only model parameters are shared among clients and the data of each client is kept private. Secondly, these algorithms rely on decision trees which may pose a security risk in a FL setting when sharing decision tree models with the server, requiring enhanced privacy and security measures. In addition, these studies primarily operate in a streaming context, with one example arriving at a time, which differs from our setup that is characterised by the simultaneous processing of multiple examples, and contrasts with the typical use of neural networks in FL.
@@ III Problem Statement
We first define the following notations used throughout the paper. We assume there exist $K$ clients and $T$ timesteps in the FL setting. In our context, a timestep refers to a specific interval during which new data is received by a client, which is composed of $R$ communication rounds. The collection of private data received by a client $k$ at timestep $t$ is represented as $D_{k}^{t}\{X,Y\},k\in K,t\in T$. Each dataset $D_{k}^{t}$ contains $|D_{k}^{t}|$ instances. We further assume a classification setting, where $X$ is the input space, $Y$ is the output space, $\hat{Y}$ represents the predicted class, and $S$ is the sensitive attribute of $X$.
@@ III-B Concept Drift
  • For group $S=s$, there is a concept drift if $P^{t}(y|X,S=s)\neq P^{t-1}(y|X,S=s)$ for any $t>1$.
  • For group $S\neq s$, there is no concept drift if $P^{t}(y|X,S\neq s)=P^{t-1}(y|X,S\neq s)$ for any $t>1$.
In other words, group-specific concept drift occurs when one group experiences a change in the conditional distribution of the target variable over time, while other groups’ conditional distribution remain constant. This difference in concept drift between groups can result in a decrease in fairness, even if loss remains fairly stable. This occurs due to changes in loss within group $S=s$ attributed to concept drift that may not be readily noticeable in the overall loss, particularly when there is an imbalance in the dataset with respect to the sensitive attribute (i.e. when group $S=s$ is underrepresented).
In these situations, conventional drift detection mechanisms that rely solely on overall loss metrics may struggle to capture these underlying dynamics. This underscores the need to incorporate group-specific loss monitoring over time, as fairness concerns can arise when shifts in concept drift affect opportunities for favorable outcomes across groups. It is important to note that, although studies have delved into fairness under concept drift in centralized environments, the concept of group-specific concept drift had not been formally addressed in the literature. In addition, empirical experiments to validate its impact are lacking as the literature focuses on general concept drift, overlooking the nuances introduced by group-specific variations.
@@
  • For each client $k$, there can be a group-specific concept drift i.e., $P^{t}_{k}(y|X,S=s)\neq P^{t-1}_{k}(y|X,S=s)$ and $P^{t}_{k}(y|X,S\neq s)=P^{t-1}_{k}(y|X,S\neq s)$ for any $t>1$.
@@
Figure 1 illustrates five different concept drift scenarios that can arise in FL with three distinct concepts (‘A’, ‘B’ and ‘C’). We focus on studying scenario 4 - group-specific distributed concept drift, as it presents a unique and unexplored aspect of concept drift in the FL setting. In this scenario it can be observed that group-specific concepts differ across time and across clients. By understanding and addressing the intricacies of this scenario, we aim to enhance the fairness of federated models when clients and sensitive attributes experience group-specific distributed concept drifts.

### v4

@@ Abstract
In the evolving field of machine learning, ensuring group fairness has become a critical concern, prompting the development of algorithms designed to mitigate bias in decision-making processes. Group fairness refers to the principle that a model’s decisions should be equitable across different groups defined by sensitive attributes such as gender or race, ensuring that individuals from privileged groups and unprivileged groups are treated fairly and receive similar outcomes. However, achieving fairness in the presence of group-specific concept drift remains an unexplored frontier, and our research represents pioneering efforts in this regard. Group-specific concept drift refers to situations where one group experiences concept drift over time while another does not, leading to a decrease in fairness even if accuracy remains fairly stable. Within the framework of Federated Learning, where clients collaboratively train models, its distributed nature further amplifies these challenges since each client can experience group-specific concept drift independently while still sharing the same underlying concept, creating a complex and dynamic environment for maintaining fairness. The most significant contribution of our research is the formalization and introduction of the problem of group-specific concept drift and its distributed counterpart, shedding light on its critical importance in the field of fairness. Additionally, leveraging insights from prior research, we adapt an existing distributed concept drift adaptation algorithm to tackle group-specific distributed concept drift which uses a multi-model approach, a local group-specific drift detection mechanism, and continuous clustering of models over time. The findings from our experiments highlight the importance of addressing group-specific concept drift and its distributed counterpart to advance fairness in machine learning.
@@
© 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. This is the author’s accepted version of the paper: T. Salazar, J. Gama, H. Araújo and P. H. Abreu, ‘‘Unveiling Group-Specific Distributed Concept Drift: A Fairness Imperative in Federated Learning’’, IEEE Transactions on Neural Networks and Learning Systems, early access, September 2025. DOI: 10.1109/TNNLS.2025.3601834
@@ I Introduction
Fairness is a fundamental concern in machine learning, aiming to prevent discrimination in algorithmic decision-making processes with respect to sensitive attributes such as race or gender [1, 2, 3]. Ensuring group fairness means that models should provide equitable outcomes across different groups defined by these sensitive attributes. For instance, the COMPAS (Correctional Offender Management Profiling for Alternative Sanctions) tool [4], used in criminal justice for assessing recidivism risk, has been criticized for exhibiting racial bias. Specifically, it was found to disproportionately label Black defendants as high risk compared to white defendants, despite having similar risk profiles. This case underscores the critical importance of addressing fairness in machine learning systems and highlights the potential consequences of biased algorithms.
However, achieving fairness becomes even more challenging in the presence of concept drift, which refers to the phenomenon where the underlying data distribution changes over time. Concept drift generally impacts overall model accuracy and performance, as models may struggle to adapt to new patterns in the data. Specifically, group-specific concept drift, a concept we introduce and formalize in this paper, occurs when the data distribution for a specific group, characterized by a sensitive attribute, changes over time, while the distribution for other groups remains constant. This discrepancy can have significant implications for fairness, as it directly affects equitable outcomes across different groups.
@@
To illustrate the importance of the problem, consider an example of group-specific concept drift in the context of loan lending. Suppose a financial institution operates in a region where historically marginalized communities have limited access to financial services. Over time, societal changes, government initiatives, and increased awareness promote financial inclusion, leading to an increase in loan applications from these previously underserved communities. In this evolving environment, group-specific concept drift occurs as the data distribution for loan applicants from marginalized communities changes. Traditional lending models may not adequately capture the creditworthiness and financial profiles of these communities due to the historical underrepresentation and lack of relevant data. If the model fails to adapt to these changes, its predictions for these communities may become less accurate, even though the overall accuracy might remain stable. This results in reduced fairness, as the model’s outcomes become inequitable across different groups. Specifically, the decrease in accuracy for the affected group leads to a decline in group fairness. Consequently, there is a risk of perpetuating bias in loan approval decisions. Addressing group-specific concept drift is relevant to ensure fair access to loans for individuals from historically marginalized communities.
Federated Learning (FL) offers a distinctive environment for studying fairness in the context of group-specific concept drift within a distributed framework. FL is a collaborative machine learning approach where multiple clients train a shared model under the coordination of a central server, while keeping the training data decentralized [5]. The continuous operation of FL systems over extended periods makes them susceptible to gradual changes such as concept drift. This operational framework is well-suited for examining fairness in the presence of group-specific distributed concept drift, as the distributed nature of FL introduces novel challenges. Specifically, group-specific concept drift may manifest in different clients at different times, complicating efforts to maintain fairness and accuracy across the entire system. As such, in this work, we also address these challenges by introducing and formalizing the concept of group-specific distributed concept drift within distributed environments.
@@
Going back to the previous example of loan lending, now in a FL setting, consider a scenario where clients representing different geographical regions experience varying rates of change in the data distribution of loan applicants. For instance, a financial institution operating in a diverse geographic landscape may have multiple branches, each serving a unique community with its distinct characteristics. In this FL framework, if one branch undergoes changes in the data distribution for loan applicants from marginalized communities, while another branch remains unaffected, the result is a manifestation of group-specific distributed concept drift. The decentralized nature of FL implies that these drifts in data distribution occur independently across different clients at various times, introducing a layer of complexity in maintaining fairness and accuracy. As the financial landscape evolves and societal changes unfold, addressing group-specific distributed concept drift becomes imperative to uphold fairness and prevent perpetuation of biases in loan approval decisions for diverse and historically underserved communities.
@@
We pioneer the introduction and formalization of the concept of group-specific concept drift and group-specific distributed concept drift, shedding light on previously unexplored aspects within the fairness landscape.
@@
The remainder of this work is structured as follows: Section II provides some background and related work on the subject, Section III describes the problem statement, Section IV presents the proposed approach, Section V describes the methodology and experimental design, Section VI presents the results, Section VII presents the discussions and possible future research, and Section VIII concludes our work.
@@ II Related Work
II-A Group Fairness
Fairness-aware machine learning has gained significant attention, and algorithms designed to promote fairness can be categorized into three groups based on the stage at which they are implemented: pre-processing, in-processing, and post-processing [1, 2, 6, 7]. The algorithm proposed in this study belongs to the in-processing category. Fairness metrics can be broadly classified into two main groups: group fairness and individual fairness. Group fairness emphasizes that the outcomes of an algorithm do not disproportionately favor or harm any particular group based on sensitive attributes such as race, gender, ethnicity, religion, or socioeconomic status. On the other hand, individual fairness aims to ensure similar predictions for individuals who share similar characteristics [1, 6]. This research specifically focuses on optimizing group fairness.
@@ II-B Fair Federated Learning
Ensuring fairness in FL is important due to the diverse and heterogeneous nature of both the data and the participants. Several types of fairness have been identified in FL [8], including group fairness [9, 10, 11, 12, 13], which ensures equitable treatment across sensitive groups; performance fairness [14, 15], which aims for uniform model performance across clients; selection fairness [16], which provides clients an equal chance to participate; and contribution fairness [17], which allocates rewards based on each client’s contribution to the global model. Group fairness in FL, which is the focus of this work, aims to mitigate bias while preserving privacy in a decentralized environment using techniques such as fair global model aggregation [9, 10, 11, 12] and local debiasing methodologies at each client [13]. Readers can refer to [8] for a comprehensive survey on this topic.
Current fair FL approaches, designed for static FL, lack mechanisms to detect and adapt to changing data distributions, rendering them unfit for addressing fairness under group-specific distributed concept drift. In such scenarios, characteristics or patterns observed in one concept may not be applicable to others, intensifying fairness challenges in FL environments. Hence, using a multi-model approach offers a promising solution to address the complexities of concept drift while preserving fairness across diverse data distributions.
@@ II-C Federated Learning under Concept Drift
Concept drift has been extensively studied in the centralized setting, and readers can refer to the surveys [18, 19] for further details. However, applying centralized algorithms to FL is often unsuitable for handling distributed concept drift, as FL must contend with heterogeneous data distributions across time and clients, all while preserving data privacy. The decentralized nature of FL, where sensitive data remains local and only model updates are shared, poses unique challenges that traditional centralized methods cannot adequately address [20, 21, 22, 23, 24, 25]. Consequently, most existing algorithms do not fully account for the distributed and privacy-preserving requirements of FL when addressing concept drift.
In the context of FL, research on concept drift is still in its early stages and most works do not account for the distributed nature of concept drift in their algorithms [20], failing to address heterogeneity across time and clients simultaneously [21, 22, 23, 24, 25]. In their recent work, Jothimurugesan et al. [20] introduce a novel approach to tackle the challenge of distributed concept drift in FL. They treat drift adaptation as a dynamic clustering problem, emphasizing the limitations of single global models when dealing with distributed drifts. To address this issue, the authors propose a multi-model approach that leverages local drift detection based on the overall loss, offering a more robust solution to adapt to evolving data distributions over time. However, it is important to note that this algorithm, referred to as FedDrift, focuses on the global loss and does not consider group-specific losses. In the context of our work, where we specifically address group-specific distributed concept drift, FedDrift’s reliance on global loss renders it ill-suited for our scenario. Group-specific concept drift may not be effectively detected by FedDrift, as it does not account for group-specific loss variations in local models that are crucial for capturing changes specific to certain groups. This limitation underscores the need for a more nuanced approach to address fairness concerns under group-specific distributed concept drift, which we elaborate on in the subsequent sections.
@@ II-D Limitations and Research Context
While certain studies have delved into fairness under concept drift in centralized streaming environments [26, 27, 28, 29], it is important to recognize their limitations when applied to fairness under concept drift in FL. These fairness-aware algorithms may not be well-suited for FL due to several reasons. Firstly, they typically require direct access to sensitive attributes, which is restricted in FL where only model parameters are shared among clients and the data of each client is kept private. Secondly, these algorithms rely on decision trees which may pose a security risk in a FL setting when sharing decision tree models with the server, requiring enhanced privacy and security measures. In addition, these studies primarily operate in a streaming context, with one example arriving at a time, which differs from our setup that is characterized by the simultaneous processing of multiple examples, and contrasts with the typical use of neural networks in FL. Similarly, centralized methods in multistream classification [30, 31, 32] predict class labels for the target stream using labels from a source stream, where concept drift can occur asynchronously. However, these methods are both fairness-unaware and designed for streaming environments, making them ill-suited for this scenario for the same reasons.
@@ III Problem Statement
We first define the notations used throughout this paper. We assume there exist $K$ clients and $T$ timesteps in the FL setting. In our context, a timestep refers to a specific interval during which new data is received by a client, which is composed of $R$ communication rounds. The collection of private data received by a client $k\in K$ at timestep $t\in T$ is represented as:
$D_{k}^{t}=\{(x_{i},y_{i})\mid x_{i}\in X,y_{i}\in Y,i=1,\dots,|D_{k}^{t}|\}$
Each dataset $D_{k}^{t}$ contains $|D_{k}^{t}|$ instances. We assume a classification setting, where $X$ is the input space, $Y$ is the output space, $\hat{Y}$ denotes the predicted class, and $S$ represents a sensitive attribute of $X$. The sensitive attribute $S$ takes values $s\in S$, where $S$ is the set of possible values corresponding to different groups.
@@ III-B Concept Drift
  • For a specific group $S=s$, there is a concept drift if $P^{t}(y|X,S=s)\neq P^{t-1}(y|X,S=s)$ for any $t>0$.
  • For other groups $S\neq s$, there is no concept drift if $P^{t}(y|X,S\neq s)=P^{t-1}(y|X,S\neq s)$ for any $t>0$.
In other words, group-specific concept drift occurs when one group experiences a change in the conditional distribution of the target variable over time, while other groups’ conditional distribution remain constant. This difference in concept drift between groups can result in a decrease in fairness, even if loss remains fairly stable. This occurs due to changes in loss within group $s$ attributed to concept drift that may not be readily noticeable in the overall loss, particularly when there is an imbalance in the dataset with respect to the sensitive attribute (i.e. when group $s$ is underrepresented).
In these situations, conventional drift detection mechanisms that rely solely on overall loss metrics may struggle to capture these underlying dynamics. This underscores the need to incorporate group-specific loss monitoring over time, as fairness concerns can arise when concept drifts affect opportunities for favorable outcomes across groups. It is important to note that, although studies have delved into fairness under concept drift in centralized environments, the concept of group-specific concept drift had not been formally addressed in the literature. In addition, empirical experiments to validate its impact are lacking as the literature focuses on general concept drift, overlooking the nuances introduced by group-specific variations.
@@
  • For each client $k$, there can be a group-specific concept drift i.e., $P^{t}_{k}(y|X,S=s)\neq P^{t-1}_{k}(y|X,S=s)$ and $P^{t}_{k}(y|X,S\neq s)=P^{t-1}_{k}(y|X,S\neq s)$ for any $t>0$.
@@
Figure 1 illustrates five different concept drift scenarios that can arise in FL with three distinct concepts (‘A’, ‘B’ and ‘C’).
  • Scenario 0: no concept drift - the underlying data distribution remains stable over time across all clients. There is no concept drift experienced by any client, and the system operates under static conditions.
  • Scenario 1: general concept drift - all clients experience a uniform concept drift in the data distribution over time. The drift is consistent, affecting all clients equally, and reflects a global change in the overall data environment.
  • Scenario 2: distributed concept drift - each client experiences independent changes in their local data distribution over time. The drifts are uncoordinated, meaning different clients may encounter different concepts, but these drifts happen independently across clients.
  • Scenario 3: group-specific general concept drift - groups within the clients experience concept drift, but this drift happens uniformly across all clients for that group. For example, all clients experience the same concept drift for a specific sensitive group, though the drift remains general across the federation.
  • Scenario 4: group-specific distributed concept drift - groups within different clients experience concept drift at different times. The group-specific concept drifts vary not only over time but also across different clients. This scenario introduces a unique complexity, as the concept drift occurs in a distributed manner, affecting different groups within different clients at different times.
We focus on studying scenario 4 - group-specific distributed concept drift, as it presents a unique and unexplored aspect of concept drift in the FL setting. This scenario demands novel techniques that consider the granularity of group-specific variations across different clients, which cannot be effectively handled by conventional methods aimed at distributed concept drift. By addressing this gap, our work introduces a framework that ensures fairness by detecting and adapting to these group-specific concept drifts.